# Online Navigation Refinement: Achieving Lane-Level Guidance by Associating Standard-Definition and Online Perception Maps

### v3

@@ 1 Introduction
Lane-level navigation has emerged as a critical capability in geographic information systems (GIS) hansson2020lane; guo2025lane and navigation-based tasks peng2025navigscene, offering finer-grained guidance than traditional road-level navigation based on standard definition (SD) maps (Fig. 1 (a)) zhang2024survey. However, today’s lane-level navigation systems mostly rely on pre-built HD maps, which are expensive to create and maintain. They often miss real-world changes elghazaly2023high—like construction, rule updates, or accidents—leading to outdated guidance and safety risks. Recently, online perception (OP) maps have become research hotspots within autonomous driving li2022hdmapnet; maptrv2. Powered by vehicle sensors, the OP maps offer current and localized lane depictions. However, as shown in Fig. 1 (b), they inherently lack global route topology and cannot support lane-level navigation wong2020mapping.
@@
To achieve up-to-date and low-cost lane-level navigation, we introduce a new mission called Online Navigation Refinement (ONR). The core objective of the mission is to transform the road-level navigation derived from SD maps into precise lane-level navigation aligned with the OP maps. For a fast and accurate process, ONR needs a paradigm of map-to-map matching, called map association, as shown in Fig 1 (c). Compared with map matching ($M$) that bind GPS trajectories to static SD/HD maps chao2020survey as a path-to-map matching based on the HMM DBLP:conf/gis/NewsonK09, seq-to-seq model feng2020deepmm; ren2021mtrajrec, graph model liu2023graphmm and transformers tang2025elevation, map association uses map-to-map matching because SD maps and OP maps are heterogeneous and do not share a one-to-one correspondence. Attempting to directly align SD paths with OP maps, or the other way around, overlooks the semantic disparities between the two.
@@
Specifically, map association faces two key challenges. First, no public datasets provide structured, lane-to-road-level correspondences between SD maps and online perception (OP) maps. Although there exist auto-driving datasets (e.g., nuScenes caesar2020nuscenes, OpenLaneV2 wang2023openlanev2) offer local lane geometries, and OpenStreetMap provides road-level topologies, none establish explicit, learnable mappings between them. Secondly, SD and OP maps display inherent heterogeneity due to diverse granularity, resulting in substantial spatial and semantic differences. A robust association mechanism is required to manage spatial fluctuations (such as GPS drift and changes in scale) and semantic disparities (such as the differing number of lanes on a road). It should also be capable of processing noise issues in OP maps in real time, which include lane discontinuities, omissions, or errors in OP mapping. Figure 6 in Appendix D illustrates the complexity and challenge of map association.
@@
To address these challenges, we make three core contributions: (1) Dataset: We introduce Online Map Association Dataset (OMA), the first open source benchmark for the online navigation refinement mission in the paradigm of map association, derived from nuScenes caesar2020nuscenes and OpenStreetMap osm, which contains more than 30K scenarios, 480K road paths and 2.6M lane vectors with manually annotated associations. (2) Baseline: We present MAT, a lightweight transformer-based model for the real-time association of maps. To handle spatial and semantic differences, MAT incorporates two key modules: path-aware attention and spatial attention. Path-aware attention reorders and group vector tokens by path index, implicitly encoding topological structure to align SD and OP maps despite spatial offsets. Spatial attention sorts and groups road/centerline tokens using a spatial curve, enabling global context modeling through cross-topological feature integration. (3) Metric: We propose Navigation Refinement P-R (NR P-R), a metric that measures alignment of the path and the correspondence through precision, recall and F1 specifically designed for map association. Traditional $M$ metrics only measure the accuracy of the matching, but the NR P-R evaluates both the geometric similarity and the accuracy of the matching. Moreover, NR P-R only utilizes annotation in ground-truth perception maps, rendering the benchmark suitable for the evaluation of any map generation method.
@@ 2 Related Work
Map Matching. Map matching investigates how to link GPS tracks to SD or HD maps. hansson2020lane utilize a hidden Markov model (HMM) considering trajectories as observations and the corresponding road segments as states, demonstrating superiority over earlier geometric matching algorithms. In the realm of deep learning, the main approach involves sequence-to-sequence (Seq2Seq) methods feng2020deepmm; ren2021mtrajrec. With an input trajectory consisting of a sequence of geo-points, the encoder-decoder framework provides a sequence of matched road segments. Recently, GraphMM liu2023graphmm has used a graph-based approach that explicitly integrates all the correlations mentioned above, while EAM3 tang2025elevation use a BERT-like transformer alaparthi2020bidirectional to achieve the best performance. Compared to the map matching method, MAT uses a map-to-map paradigm to accommodate the many-to-one correspondence between lanes and roads in real time. Moreover, unlike the self-attention used in EAM3, MAT employs spatial and path-aware attention by incorporating the spatial curve and the path index, thus striking a balance between performance and latency.
Map Generation and Priors. The construction of online perception maps is a popular topic in autonomous driving hao2024mapdistill; zhang2024sparse; li2024mlp and is crucial for subsequent tasks wan2024sp; zhang2024p2ftrack. HDMapNet li2022hdmapnet pioneered BEV-based map generation through sensor fusion, while LSS philion2020lift introduced depth-aware BEV transformation. VectorMapNet liu2023vectormapnet enabled end-to-end vector prediction, and the MapTR liao2023maptr series introduced hierarchical query embeddings for instance-level construction. We adapt MapTRv2 maptrv2 for online HD construction using the annotations of our dataset and demonstrate its compatibility with the SD-HD association. Recent works such as SMERF luo2024augmenting, P-MapNet jiang2024p, and TopoSD yang2024toposd leverage SD map priors to mitigate sensor noise and refine perception geometry. Unlike these approaches that focus on enhancing generation quality, MAT targets the explicit association phase, utilizing topological optimization to resolve assignment ambiguity arising from residual noise.
@@ 4.1 Dataset Overview
The source of ground-truth OP maps is nuScenes nuscenes2019, which includes locations in Boston and Singapore, featuring centerline geometries scanned with LiDAR. The SD maps were obtained from OpenStreetMap (OSM). We manually annotated the data to establish the association between the ground truth of the OP map and the SD map. The visualization is shown in Fig. 3 (a), which shows the global SD and ground-truth OP map of Boston with an association annotation.
@@ 4.2 Evaluation Metric
In map matching, evaluation typically involves metrics such as Trajectory-Level Accuracy and the Ratio of Longest Common Subsequence, which presume a one-to-one mapping between trajectories and paths, concentrating only on the accuracy and sequence of path matching. Conversely, in map association, to assess various map generation methods, our goal is to use the association annotations of the ground-truth OP map to evaluate the results of the various OP maps with noise. Drawing inspiration from Reachability Precision Recall (P-R) lu2023translating, we developed the Navigation Refinement P-R metric. This allows for the assessment of any map generation method’s perception map performance for the map association task, relying solely on ground-truth OP map annotations.
@@
1. Path Matching between Prediction and Ground Truth: In the first stage, a connection is drawn between the predicted paths and GT paths within the OP map. Notably, only GT paths come with association annotations, making this matching essential for the following evaluations. A match is deemed successful if the bidirectional Chamfer distance is less than the distance threshold of 1 meter, consistent with Reachability P-R lu2023translating. Predictions that do not meet this threshold are labeled FP, thereby excluding them from further SD-HD analysis. As illustrated in Row 1 of Fig. 4, Case 1 (left) shows that the prediction with minimal deviation successfully aligns with the GT paths. While Case 1 (middle) represents the heavily-deviated prediction results in an FP label.
2. Label Sequence Alignment Verification: Successful matches are converted into simplified label sequences, indicating SD map link traversals. Failures in sequence alignment at this point lead to FP/FN labels. As shown in Case 2 (middle) of Fig. 4, the case is a misaligned sequence result which is an FP, while Case 2 (right) shows that the GT without a match leads to a false negative (FN).
3. Length-Aware Overlap Assessment: For aligned sequences, we quantify the semantic consistency by computing the Label Overlap score. Specifically, this score is defined as the ratio of the length of correctly matched segments (where the predicted road ID matches the ground truth road ID) to the total path length. We compare the length proportions of road labels between the predicted and GT sequences to determine these matches. The true positive (TP) classification is determined based on thresholds. As shown in Case 3 (left) of Fig. 4, the case is a 67% overlap, which is TP at T = 50% but FP at T=95%.
Following the mAP conventions lin2014microsoft, we use $T=[0.5:0.05:0.95]$ (10 thresholds) and report mean P-R and F1 scores. To mitigate path-length bias, we separately compute metrics across 15 length intervals $L=[[0,5),[5,10),..,[70,+\infty)]$ before aggregation. The ablation study of distance threshold and length intervals is shown in the appendix G.
@@ 5.3 Spatial Attention
2) Serialization via Space-Filling Curves: We employ a space-filling curve function $\varphi^{-1}$ (e.g., the Hilbert curve hilbert1935stetige) to map the 3D coordinates to a single 1D index. This mapping is crucial as it preserves spatial locality in the 1D domain better than simple row-major scanning.
@@ 5.4 Association and Loss Function
Association. The association between the roads and the centerline is calculated through a mechanism of cross-attention. For each road $j$, we first aggregate its token features $\{F^{road}_{j1},\ldots,F^{road}_{jN}\}$ into a representative feature $\bar{F}^{road}_{j}=\frac{1}{N}\sum_{n=1}^{N}F^{road}_{jn}$, where NN denotes the number of road tokens on roads $r_{j}$ and $F^{road}_{jn}\in\mathbb{R}^{d}$. The association probability $Prob_{ij}$ between the centerline $i$ and the road $j$ is then calculated as:
@@
Loss Function. Followed by liu2023graphmm; ren2021mtrajrec, we optimize the model using a combination of cross-entropy loss (CE) and connection temporal classification (CTC) loss. The total loss is a weighted sum:
@@ 6 Experiment
HMM DBLP:conf/gis/NewsonK09 SIGSPATIAL’09 $M$ P-M 70.1 465
DeepMM feng2020deepmm TMC’20 $M$ P-M 70.6 328
MTrajRec ren2021mtrajrec KDD’21 $M$ P-M 71.2 849
GraphMM liu2023graphmm TKDE’23 $M$ P-M 72.3 469
EAM3 tang2025elevation TITS’25 $M$ P-M 72.9 345
KNN stuart2010matching GM M-M 68.6 299
GMT he2024learnable TPAMI’24 GM M-M 68.8 93
FastMAC zhang2024fastmac CVPR‘24 PM M-M 70.3 62
@@
Table 2: Result on OMA Test set. La. means latency. $M$, PM, GM, MA means map matching, graph matching, point matching and map association method. P-M and M-M means the model is trained and inferred by path-to-map or map-to-map. SGG. means SeqGrowGrpah xie2025seqgrowgraph.
@@
HMM DBLP:conf/gis/NewsonK09 SIGSPATIAL’09 $M$ P-M 33.1 36.0 46.5 561
DeepMM feng2020deepmm TMC’20 $M$ P-M 32.8 33.5 43.9 733
MTrajRec ren2021mtrajrec KDD’21 $M$ P-M 34.1 36.8 48.9 1593
GraphMM liu2023graphmm TKDE’23 $M$ P-M 35.9 38.7 49.2 889
EAM3 tang2025elevation TITS’25 $M$ P-M 36.3 39.1 50.8 679
KNN stuart2010matching GM M-M 31.7 34.6 43.5 313
GMT he2024learnable TPAMI’24 GM M-M 30.1 33.4 42.3 105
FastMAC zhang2024fastmac CVPR‘24 PM M-M 32.0 35.8 45.0 81
@@ 6.2 Result
To evaluate the performance of different methods on the OMA dataset, we categorize existing approaches into map matching methods DBLP:conf/gis/NewsonK09; feng2020deepmm; ren2021mtrajrec; liu2023graphmm; tang2025elevation, graph matching methods he2024learnable, and point matching methods zhang2024fastmac. For map matching, we traverse every path on the OP map as a GPS trajectory to match against the SD map. For graph and point matching, we treat the SD and OP maps as separate graphs or point clouds.
Val set. As shown in Table 1, MAT achieves the optimal balance between precision and efficiency on the validation set. Specifically, compared to traditional, Seq2Seq, and graph-based map matching methods, MAT-T demonstrates significant improvements in $\text{NR-F1}^{50:95}$, outperforming HMM DBLP:conf/gis/NewsonK09 by 8.1%, DeepMM feng2020deepmm by 7.6%, MTrajRec ren2021mtrajrec by 7.0%, and the recent EAM3 tang2025elevation by 5.3%, all while maintaining a low inference latency of 34 ms. Moreover, MAT-T achieves superior performance compared to graph matching techniques like GMT he2024learnable and point matching approaches such as FastMAC zhang2024fastmac, proving its effectiveness in handling the heterogeneity between SD and OP maps which typically challenges traditional matching or point cloud algorithms.
Test set. To further verify the model’s robustness against varying noise patterns crucial for real-world deployment, we extended the evaluation on the Test set to include OP maps generated by three distinct methods: MapTR liaomaptr, MapTRv2 maptrv2, and SeqGrowGraph xie2025seqgrowgraph, as summarized in Table 2. The results demonstrate that our proposed MAT method consistently outperforms all state-of-the-art baselines across these disparate generators without requiring specific fine-tuning. On the SeqGrowGraph part, MAT-L achieves an $\text{NR-F1}^{50:95}$ of 54.9%, surpassing the strongest baseline EAM3 tang2025elevation by 4.1%, while on the more challenging MapTR dataset, which exhibits different topological error patterns, MAT-L maintains a significant lead of 5.6% over EAM3. This consistent superiority confirms that MAT’s architecture effectively generalizes across distinct noise distributions—ranging from fragmentation in MapTR to connectivity issues in growing-based methods—thereby validating the strong cross-generator generalization ability of the proposed Online Navigation Refinement framework.

### v4

@@ 1 Introduction
Lane-level navigation has emerged as a critical capability in geographic information systems (GIS) Hansson et al. (2020); Guo et al. (2025) and navigation-based tasks Peng et al. (2025); Li et al. (2025); Shan et al. (2025), offering finer-grained guidance than traditional road-level navigation based on standard definition (SD) maps (Fig. 1 (a)) Zhang et al. (2024d). However, today’s lane-level navigation systems mostly rely on pre-built HD maps, which are expensive to create and maintain. They often miss real-world changes Elghazaly et al. (2023)—like construction, rule updates, or accidents—leading to outdated guidance and safety risks. Recently, online perception (OP) maps have become research hotspots within autonomous driving Li et al. (2022); Liao et al. (2023b). Powered by vehicle sensors, the OP maps offer current and localized lane depictions. However, as shown in Fig. 1 (b), they inherently lack global route topology and cannot support lane-level navigation Wong et al. (2020).
@@
To achieve up-to-date and low-cost lane-level navigation, we introduce a new mission called Online Navigation Refinement (ONR). The core objective of the mission is to transform the road-level navigation derived from SD maps into precise lane-level navigation aligned with the OP maps. For a fast and accurate process, ONR needs a paradigm of map-to-map matching, called map association, as shown in Fig 1 (c). Compared with map matching (MM) that bind GPS trajectories to static SD/HD maps Chao et al. (2020) as a path-to-map matching based on the HMM Newson (2009), seq-to-seq model Feng et al. (2020); Ren et al. (2021), graph model Liu et al. (2023b) and transformers Tang et al. (2025), map association uses map-to-map matching because SD maps and OP maps are heterogeneous and do not share a one-to-one correspondence. Attempting to directly align SD paths with OP maps, or the other way around, overlooks the semantic disparities between the two.
@@
Specifically, map association faces two key challenges. First, no public datasets provide structured, lane-to-road-level correspondences between SD maps and online perception (OP) maps. Although there exist auto-driving datasets (e.g., nuScenes Caesar et al. (2020a), OpenLaneV2 Wang et al. (2023)) offer local lane geometries, and OpenStreetMap provides road-level topologies, none establish explicit, learnable mappings between them. Secondly, SD and OP maps display inherent heterogeneity due to diverse granularity, resulting in substantial spatial and semantic differences. A robust association mechanism is required to manage spatial fluctuations (such as GPS drift and changes in scale) and semantic disparities (such as the differing number of lanes on a road). It should also be capable of processing noise issues in OP maps in real time, which include lane discontinuities, omissions, or errors in OP mapping. Figure 6 in Appendix D illustrates the complexity and challenge of map association.
@@
To address these challenges, we make three core contributions: (1) Dataset: We introduce Online Map Association Dataset (OMA), the first open source benchmark for the online navigation refinement mission in the paradigm of map association, derived from nuScenes Caesar et al. (2020a) and OpenStreetMap 31, which contains more than 30K scenarios, 480K road paths and 2.6M lane vectors with manually annotated associations. (2) Baseline: We present MAT, a lightweight transformer-based model for the real-time association of maps. To handle spatial and semantic differences, MAT incorporates two key modules: path-aware attention and spatial attention. Path-aware attention reorders and group vector tokens by path index, implicitly encoding topological structure to align SD and OP maps despite spatial offsets. Spatial attention sorts and groups road/centerline tokens using a spatial curve, enabling global context modeling through cross-topological feature integration. (3) Metric: We propose Navigation Refinement P-R (NR P-R), a metric that measures alignment of the path and the correspondence through precision, recall and F1 specifically designed for map association. Traditional MM metrics only measure the accuracy of the matching, but the NR P-R evaluates both the geometric similarity and the accuracy of the matching. Moreover, NR P-R only utilizes annotation in ground-truth perception maps, rendering the benchmark suitable for the evaluation of any map generation method.
@@ 2 Related Work
Map Matching. Map matching investigates how to link GPS tracks to SD or HD maps. Hansson et al. (2020) utilizes a hidden Markov model (HMM) considering trajectories as observations and the corresponding road segments as states, demonstrating superiority over earlier geometric matching algorithms. In the realm of deep learning, the main approach involves sequence-to-sequence (Seq2Seq) methods Feng et al. (2020); Ren et al. (2021). With an input trajectory consisting of a sequence of geo-points, the encoder-decoder framework provides a sequence of matched road segments. Recently, GraphMM Liu et al. (2023b) has used a graph-based approach that explicitly integrates all the correlations mentioned above, while EAM3 Tang et al. (2025) uses a BERT-like transformer Alaparthi and Mishra (2020) to achieve the best performance. Compared to the map matching method, MAT uses a map-to-map paradigm to accommodate the many-to-one correspondence between lanes and roads in real time. Moreover, unlike the self-attention used in EAM3, MAT employs spatial and path-aware attention by incorporating the spatial curve and the path index, thus striking a balance between performance and latency.
Map Generation and Priors. The construction of online perception maps is a popular topic in autonomous driving Hao et al. (2024); Zhang et al. (2024a); Li et al. (2024) and is crucial for subsequent tasks Wan et al. (2025); Zhang et al. (2024b). HDMapNet Li et al. (2022) pioneered BEV-based map generation through sensor fusion, while LSS Philion and Fidler (2020) introduced depth-aware BEV transformation. VectorMapNet Liu et al. (2023a) enabled end-to-end vector prediction, and the MapTR Liao et al. (2023a) series introduced hierarchical query embeddings for instance-level construction. We adapt MapTRv2 Liao et al. (2023b) for online HD construction using the annotations from our dataset and demonstrate its compatibility with the SD-HD association. Recent works such as SMERF Luo et al. (2024), P-MapNet Jiang et al. (2024), and TopoSD Yang et al. (2024) leverage SD map priors to mitigate sensor noise and refine perception geometry. Unlike these approaches that focus on enhancing generation quality, MAT targets the explicit association phase, utilizing topological optimization to resolve assignment ambiguity arising from residual noise.
@@ 4.1 Dataset Overview
The source of ground-truth OP maps is nuScenes Caesar et al. (2020b), which includes locations in Boston and Singapore, featuring centerline geometries scanned with LiDAR. The SD maps were obtained from OpenStreetMap (OSM). We manually annotated the data to establish the association between the ground truth of the OP map and the SD map. The visualization is shown in Fig. 3 (a), which shows the global SD and ground-truth OP map of Boston with an association annotation.
@@ 4.2 Evaluation Metric
In map matching, evaluation typically involves metrics such as Trajectory-Level Accuracy and the Ratio of Longest Common Subsequence, which presume a one-to-one mapping between trajectories and paths, concentrating only on the accuracy and sequence of path matching. Conversely, in map association, to assess various map generation methods, our goal is to use the association annotations of the ground-truth OP map to evaluate the results of the various OP maps with noise. Drawing inspiration from Reachability Precision Recall (P-R) Lu et al. (2023), we developed the Navigation Refinement P-R metric. This allows for the assessment of any map generation method’s perception map performance for the map association task, relying solely on ground-truth OP map annotations.
@@
1. Path Matching: Predicted paths are matched to GT via a 1m bidirectional Chamfer distance threshold Lu et al. (2023). This step is vital as only GT paths contain association annotations. Non-matches are labeled FP and excluded from SD-HD analysis. As in Fig. 4 (Row 1), low-deviation paths align (left), while high-deviation ones yield FPs (middle).
2. Label Sequence Alignment: Successful matches are converted into SD map link sequences. Alignment failures are marked FP/FN. Case 2 in Fig. 4 shows sequence misalignment leading to an FP (middle) and an unmatched GT as an FN (right).
3. Length-Aware Overlap: Semantic consistency is measured by a Label Overlap score—the length ratio of road-ID-consistent segments to the total path. TP classification depends on threshold $T$. As shown in Fig. 4 (Case 3), 67% overlap is a TP at $T=50\%$ but an FP at $T=95\%$.
Following the mAP conventions Lin et al. (2014), we use $T=[0.5:0.05:0.95]$ (10 thresholds) and report mean P-R and F1 scores. To mitigate path-length bias, we separately compute metrics across 15 length intervals $L=[[0,5),[5,10),..,[70,+\infty)]$ before aggregation. The ablation study of distance threshold and length intervals is shown in the appendix G.
@@ 5.3 Spatial Attention
2) Serialization via Space-Filling Curves: We employ a space-filling curve function $\varphi^{-1}$ (e.g., the Hilbert curve Hilbert (1935)) to map the 3D coordinates to a single 1D index. This mapping is crucial as it preserves spatial locality in the 1D domain better than simple row-major scanning.
@@ 5.4 Association and Loss Function
Association. The association between the roads and the centerline is calculated through a mechanism of cross-attention. For each road $j$, we first aggregate its token features $\{F^{road}_{j1},\ldots,F^{road}_{jN}\}$ into a representative feature $\bar{F}^{road}_{j}=\frac{1}{N}\sum_{n=1}^{N}F^{road}_{jn}$, where $N$ denotes the number of road tokens on roads $r_{j}$ and $F^{road}_{jn}\in\mathbb{R}^{d}$. The association probability $Prob_{ij}$ between the centerline $i$ and the road $j$ is then calculated as:
@@
Loss Function. Followed by Liu et al. (2023b); Ren et al. (2021), we optimize the model using a combination of cross-entropy loss (CE) and connection temporal classification (CTC) loss. The total loss is a weighted sum:
@@ 6 Experiment
HMM Newson (2009) SIGSPATIAL’09 $M$ P-M 70.1 465
DeepMM Feng et al. (2020) TMC’20 $M$ P-M 70.6 328
MTrajRec Ren et al. (2021) KDD’21 $M$ P-M 71.2 849
GraphMM Liu et al. (2023b) TKDE’23 $M$ P-M 72.3 469
EAM3 Tang et al. (2025) TITS’25 $M$ P-M 72.9 345
KNN Cochran (1965) GM M-M 68.6 299
GMT He et al. (2024) TPAMI’24 GM M-M 68.8 93
FastMAC Zhang et al. (2024e) CVPR‘24 PM M-M 70.3 62
@@
Table 2: Result on OMA Test set. La. means latency. $M$, PM, GM, MA means map matching, graph matching, point matching and map association method. P-M and M-M means the model is trained and inferred by path-to-map or map-to-map. SGG. means SeqGrowGrpah Xie et al. (2025a).
@@
HMM Newson (2009) SIGSPATIAL’09 $M$ P-M 33.1 36.0 46.5 561
DeepMM Feng et al. (2020) TMC’20 $M$ P-M 32.8 33.5 43.9 733
MTrajRec Ren et al. (2021) KDD’21 $M$ P-M 34.1 36.8 48.9 1593
GraphMM Liu et al. (2023b) TKDE’23 $M$ P-M 35.9 38.7 49.2 889
EAM3 Tang et al. (2025) TITS’25 $M$ P-M 36.3 39.1 50.8 679
KNN Cochran (1965) GM M-M 31.7 34.6 43.5 313
GMT He et al. (2024) TPAMI’24 GM M-M 30.1 33.4 42.3 105
FastMAC Zhang et al. (2024e) CVPR‘24 PM M-M 32.0 35.8 45.0 81
@@ 6.2 Result
To evaluate the performance of different methods on the OMA dataset, we categorize existing approaches into map matching methods Newson (2009); Feng et al. (2020); Ren et al. (2021); Liu et al. (2023b); Tang et al. (2025), graph matching methods He et al. (2024), and point matching methods Zhang et al. (2024e). For map matching, we traverse every path on the OP map as a GPS trajectory to match against the SD map. For graph and point matching, we treat the SD and OP maps as separate graphs or point clouds.
Val set. As shown in Table 1, MAT achieves the optimal balance between precision and efficiency on the validation set. Specifically, compared to traditional, Seq2Seq, and graph-based map matching methods, MAT-T demonstrates significant improvements in $\text{NR-F1}^{50:95}$, outperforming HMM Newson (2009) by 8.1%, DeepMM Feng et al. (2020) by 7.6%, MTrajRec Ren et al. (2021) by 7.0%, and the recent EAM3 Tang et al. (2025) by 5.3%, all while maintaining a low inference latency of 34 ms. Moreover, MAT-T achieves superior performance compared to graph matching techniques like GMT He et al. (2024) and point matching approaches such as FastMAC Zhang et al. (2024e), proving its effectiveness in handling the heterogeneity between SD and OP maps which typically challenges traditional matching or point cloud algorithms.
Test set. To further verify the model’s robustness against varying noise patterns crucial for real-world deployment, we extended the evaluation on the Test set to include OP maps generated by three distinct methods: MapTR Liao et al. , MapTRv2 Liao et al. (2023b), and SeqGrowGraph Xie et al. (2025a), as summarized in Table 2. The results demonstrate that our proposed MAT method consistently outperforms all state-of-the-art baselines across these disparate generators without requiring specific fine-tuning. On the SeqGrowGraph part, MAT-L achieves an $\text{NR-F1}^{50:95}$ of 54.9%, surpassing the strongest baseline EAM3 Tang et al. (2025) by 4.1%, while on the more challenging MapTR dataset, which exhibits different topological error patterns, MAT-L maintains a significant lead of 5.6% over EAM3. This consistent superiority confirms that MAT’s architecture effectively generalizes across distinct noise distributions—ranging from fragmentation in MapTR to connectivity issues in growing-based methods—thereby validating the strong cross-generator generalization ability of the proposed Online Navigation Refinement framework.